OpenStudy (anonymous):

this a question on behalf of@PeepzStudyTime please answer so I can give her the help she would like ABCD is a rectangle with diagonals AC and BD intersect at E. If AE=x+2 and BD= 4x-16. Find the length of the diagonals of rectangle? Please reply SOON @igreen please help

OpenStudy (igreen):

Is there a picture? @PeepzStudyTime

OpenStudy (anonymous):

@lloio thnx :) @iGreen Nope theres no pic

OpenStudy (igreen):

Okay, well the square's gonna look something like:

OpenStudy (anonymous):

@iGreen its a rectangle :(

OpenStudy (anonymous):

@iGreen not a square but anyway how to do it

OpenStudy (igreen):

xD It's okay.

OpenStudy (anonymous):

@iGreen XP

OpenStudy (igreen):

Okay, these are congruent so we can write: \(4x - 16 = 2(x + 2)\) Distribute the 2 into the parenthesis: \(4x - 16 = 2x + 4\) Add 16 to both sides: \(4x = 2x + 20\) Subtract 2x to both sides: \(2x = 20\) Divide 2 to both sides: Can you finish it?

OpenStudy (igreen):

Well it's going to be: \(x = 10\)

OpenStudy (igreen):

We can plug this back into the equation: \(2x + 4\) \(2(10) + 4\) \(20 + 4\) \(24\) So the whole length is 24.

OpenStudy (anonymous):

thanks so much

OpenStudy (igreen):

No problem!

OpenStudy (anonymous):

is this also you @greenlegodude57

OpenStudy (anonymous):

@igreen

OpenStudy (igreen):

Yep. I kept getting banned..someone kept reporting me.

OpenStudy (anonymous):

oh ok I have him fanned to

OpenStudy (anonymous):

@lloio @iGreen Thnx :)

OpenStudy (igreen):

If you get reported a lot of times, then you get banned automatically.

OpenStudy (anonymous):

so welcome

OpenStudy (igreen):

So I had to switch to my other account temporarily.

OpenStudy (anonymous):

oh so you say to have two accounts just in case

OpenStudy (anonymous):

bye

OpenStudy (igreen):

Well, not really..my first account was 'greenlegodude57', I just made a new one 'cause I didn't like the name..but yeah I use it as a spare.